OpenStudy (anonymous):

need help!!! use subtraction theorm to write equivalent expressions : 8y-9x

OpenStudy (anonymous):

What is the subtraction theorem? I mean, just look at the expression, there isnt anything that can be done. They arent like terms so they cant be subtracted.

OpenStudy (anonymous):

just looking*

OpenStudy (anonymous):

so what should i put??

OpenStudy (anonymous):

I dont know >.< there are a million different things that are equivalent. If i had to take a guess, it would be one of these two: \[-9x+8y\]All i did here was switch the order. Or maybe they want: \[8y+(-9x)\] Maybe that is the "subtraction theorem"? That:\[a-b=a+(-b)\]

OpenStudy (anonymous):

Whoever is telling you its called the "subtraction theorem", tell them to stop making up fancy names lol <.<

OpenStudy (anonymous):

lol its in the algebra 2 book..

OpenStudy (anonymous):

Who wrote that book? im gonna send an angry email...

OpenStudy (anonymous):

lolol, jk jk.

OpenStudy (anonymous):

haha thanks for the help though i dont kno how to do any of these problems at all lol..

OpenStudy (anonymous):

\[3\left| x+2 \right| for x=-4\] need help how do u do this!

OpenStudy (anonymous):

so the problem is asking what is the value of: \[3|x+2|\] when x = 4?

OpenStudy (anonymous):

X=-4

OpenStudy (anonymous):

whoops! my bad, i see it now. ok, so we are going to let x = -4. So everywhere we see an 'x' in the math expression, we change it to -4. that gives: \[3|x+2| \iff 3|-4+2|\]make sense so far?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

ok, so now we have to actually evaluate it. Those vertical bars "|", are called absolute value bars, and we treat them like parenthesis, so we will calculate whatever is inside them first. So, we just need to find out what -4+2 is.

OpenStudy (anonymous):

an thats -2

OpenStudy (anonymous):

right :) so now our problem is: \[3|-4+2| = 3|-2|\] Next we need to calculate what the absolute value of -2 is. There is a pretty easy rule for getting the absolute value of a number. You just make it positive. If its already positive, leave it alone. Here are some examples: |-10| = 10, |10| = 10, |-12341234| = 12341234, |-whatever| = whatever So what would |-2| be?

OpenStudy (anonymous):

2

OpenStudy (anonymous):

right again. so now we have: \[3|-2| = 3(2)\] now we just multiply to get our final answer. \[3(2) = 6\] So, thats the final answer. When x = -4, 3|x+2| = 6

OpenStudy (anonymous):

thank you so much! i get it now. so for \[\left| p+q \right|\] for p=21 and q= -9 you do \[\left| 21+-9 \right| or \left| 21-9 \right|\]?